# Efficient Exploration and Discriminative World Model Learning with an Object-Centric Abstraction

### v1

@@ 6 Discussion
Limitations The major limitation of the work is that the construction of Ab-MDP requires the existence of an object-centric mapping. While we show how these components can be learned in Section 4.3, learning the object mapping still requires a labelled dataset. Promisingly, the field as a whole is moving in a direction where abstractions are becoming more easily obtainable. For instance, object-centric abstract maps can be acquired using segmentation models (Kirillov et al.,, 2023) and visual language models (Alayrac et al.,, 2022). Similarly, the short-horizon (low level) manipulation policies can be acquired more efficiently through imitation learning. In the robotics setting it’s even been found that using off-the-shelf models for these perception and skill policies is possible (Liu et al.,, 2024). Nevertheless, a method to automatically discover object-centric representations remains an open question.
7 Acknowledgements
This work is supported by ONR MURI #N00014-22-1-2773 and ONR #N00014-21-1-2758. AGC is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), PGSD3-559278-2021. Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), PGSD3-559278-2021. We are grateful for insightful discussions with Rajesh Ranganath. We thank Ben Evans, Siddhant Haldar, Ulyana Piterbarg, Dongyan Lin, and Mandana Samiei for their helpful feedback on earlier drafts of this paper.
Appendix
@@ B.1 Abstracted MDP
Here we re-iterate our framework and discuss it in greater details. We first define a low level (reward-free) Markov Decision Process (MDP, Puterman, (1994)) as the tuple $\langle\mathcal{S},\mathcal{A},P\rangle$, with (low level) state space $\mathcal{S}$, primitive action space $\mathcal{A}$, and transition probability function $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$. Suppose the agent interacts with the environments at discrete time-steps $u=0,1,2,...$, then $P(s_{u+1}|s_{u},a_{u})$ describes the probability of transitioning to state $s_{u+1}\in\mathcal{S}$ after one time-step when choosing action $a_{u}\in\mathcal{A}$ in the current low level state $s_{u}$.
@@
B.1.1 Interpretation as Options
It is natural to interpret our hierarchical framework within the Options framework (Sutton et al.,, 1999). An option consists of a policy $\pi:\Omega\times\mathcal{A}\rightarrow[0,1]$, a termination function $\beta:\Omega\rightarrow[0,1]$, and an initiation set of states $\mathcal{I}\subseteq\mathcal{S}$. $\Omega$ is used to denote the set of all possible historical states.
@@ B.1.2 A note on semi-MDP
The decision process that selects amongst options is a semi-MDP (Theorem 1, (Sutton et al.,, 1999)). The general way to model semi-MDP dynamic requires considering the amount of low level timesteps that passes between each abstract level decision as a random variable (Sutton et al.,, 1999; Barto and Mahadevan,, 2003). Instead, we chose to simply treat the abstracted level as having a fixed time interval in between steps. This is chosen for simplicity as our focus is with learning efficiency on the abstract level, and this formulation allow us to directly apply efficient algorithms developed to solve MDPs. We also see empirically that given our $\mathcal{M}$ and $\mathcal{B}$ this choice does not lead to a less optimal policy compared to algorithms that directly solve the low level MDP.
B.1.3 Multi-Item Behaviours
@@ B.2.1 The generality of Ab-MDP
One may worry that the abstraction specified by Ab-MDP is overly specific to a particular problem and an appropriate abstraction would not exist for a different problem. To this end, we make the following observation that any sparse reward RL task can be described by a Ab-MDP with a single item,
@@ C.1 Model Fitting
Appendix C.2 contains architectural details for $f_{\theta}$. All in all, we found that our sampling method and loss function lead to a stable, simple-to-optimize objective. Moreover, it is possible to directly use $\rho$ to do planning, as is done in Zhang et al., (2018). However, we show in Section F.1.3 that it is more efficient to use our parametric modelling approach as it generalizes better to new objects.
@@ C.2 Model Architecture
We use a small key-query attention architecture that takes in sets of vectors to produce a binary prediction, see Figure 11.
@@ C.3 Exploration
Intuitively, a state with many visits (large $\text{N}(X,b)$) will have lower reward, while $T$ prevents the intrinsic reward from shrinking to zero (so the agent explores indefinitely). Indeed, in the bandits setting, this reward function is maximized when all states are visited uniformly (Zhang et al.,, 2018). We emphasize that all of our model training is done with this intrinsic reward, with no task-specific information injected at training time.
@@ Appendix D Environments
Figure 12: Example episode in an example 2D Crafting environment, figure taken from Chen et al., (2020). To solve this task, the agent needs to pickup the pickaxe, mine the cobblestone stash to get cobblestone, then use the acquired cobblestone to craft cobblestone stairs. The items must be acquired in that order as each step requires having the item from the previous step (with an additional obstacle of being able to cross the door only if it has a key).
We adopt three crafting environments proposed in Chen et al., (2020), which contain Minecraft-like crafting logic. The agent needs to traverse a grid world, collect resources, and build increasingly more complex items by collecting and crafting ingredients (e.g. getting an axe, chopping tree to get wood, before being able to make wooden planks and furniture that depend on planks). Primitive actions consist of movement in the four cardinal directions (N,E,S,W), toggle (of a switch to open doors), grab (e.g. pickaxe), mine (e.g. ores), craft (at crafting bench of specific items). The low level observation space consist of a dictionary containing: (i) an (5, 5, 300) matrix describing a $5\times 5$ world with each state being described by an 300-dim word embedding, (ii) an 300-dim inventory embedding describing things in the inventory, and (iii) a 300-dim goal embedding describing in words the task to complete. We describe each environment’s crafting sequences required to finish each task in Table 1.
@@ D.2 MiniHack
We also use the MiniHack environment (Samvelyan et al.,, 2021), which is built on the Nethack Learning Environment (Küttler et al.,, 2020). We use five hard-exploration games from Henaff et al., (2022) requiring interaction with different types of objects. The base game has a multi-modal observation space involving image, text, inventory, etc. Solving the game from the base observation is extremely difficult since each environment has multiple variants of an object (e.g. Freeze-Wand has 27 different types of wands), and randomly generated object locations. Indeed, Henaff et al., (2022) found that performant RL methods such as IMPALA (Espeholt et al.,, 2018; Küttler et al.,, 2019) and all tested global exploration methods (ICM (Pathak et al.,, 2017) and RND (Burda et al.,, 2019)) fail to solve these environments.
@@
We then embed the object identities and attributes before providing them to the agents. For object identities, we take each object’s glyph number, and map it to the NetHack text description for that glyph (which we can extract from the screen_descriptions modality in the low level observation). For each glyph’s text description (which is a short sentence), we pre-process the sentence to remove stop words, then take each word’s GloVe (Pennington et al.,, 2014) embedding and averaging over the sentence. We use the 300-dimensional Wikipedia 2014 + Gigaword 5 GloVe word vector embeddings from https://nlp.stanford.edu/projects/glove/ (6B tokens, 400k vocab, uncased). The 300-dim vectors were chosen as we found them to have naturally good separation between object types in preliminary analysis.
@@ D.2.1 Sandbox Environment
Figure 14: Sandbox Minihack environment
We design a sandbox environment where the agent can observe and interact with all items it may encounter in one of the five evaluation environments. In the sandbox, each of these objects has an 0.5 chance (independently) of spawning at a random location: levitate boots, levitate rings, levitate potions, freeze horn, freeze wands. Further there is a 0.5 change that a second of the same object type will spawn. The agent can interact with any of the objects for up to 250 low level frames, or the episode terminates once the agent sets any of the object’s attribute to “used”. Figure 14 shows an instantiation of an episode.
@@ D.2.2 Low Level Policy Errors
    To use an item in the inventory, the policy needs to select which item to use (e.g. to drink a potion the policy needs to first do primitive action "quaff", then primitive action that selects the inventory object to drink, etc.). When there are multiple objects in the inventory it just picks a random object, meaning it might not pick the object specified by the behaviour
@@
    The “go stand on” policy is coded to move in a straight line towards an object. Since we measure a policy "success" whenever the abstract state changes, if the agent walks over another object en route, the Ab-MDP will register an abstract state change and deem the transition unsuccessful
@@
    To successfully cast a freeze spell, the NetHack game needs to return a message saying the spell has bounced off a wall. To our knowledge, this is stochastic—we got as far as figuring out that the spell needs to travel sufficient distance to bounce, but even so we were unable to produce a low level policy that results in spell bounce 100% of the time. This succeeds 8̃0% of the time

### v3

@@ 6 Discussion
Limitations The major limitation of the work is that the construction of Ab-MDP requires the existence of an object-centric mapping. While we show how these components can be learned in Section 4.3, learning the object mapping still requires a labelled dataset, and a method to automatically discover object-centric representations in the form of items and attributes remains an open question. Promisingly, the field as a whole is moving in a direction where abstractions are becoming more easily obtainable. For instance, object-centric abstract maps can be acquired using segmentation models (Kirillov et al., 2023) and visual language models (Alayrac et al., 2022). Similarly, the short-horizon (low-level) manipulation policies can be acquired more efficiently through imitation learning. For example, it has been found in robotics that simply using off-the-shelf models for perception and skill policies is possible (Liu et al., 2024). Finally, we currently focus only on discrete attributes, meaning the Ab-MDP cannot precisely capture continuously varying quantities (e.g. position and velocity). However, this is only a restriction in the abstract level, and does not impose any fundamental restriction on the kind of learnable policies in the base MDP. In other words, Ab-MDP models discrete, abstract relationships present in the base MDP, while low-level policies remain free to utilize continuous features and output continuous actions in the base MDP to drive item-attribute changes for the Ab-MDP.
Overall, as foundational models advance, the field is progressively moving toward working with discrete, semantically meaningful representations. Concurrently, reinforcement learning provides rigorous and principled frameworks for designing intelligent agents, encompassing exploration, planning, and hierarchical behaviours. We hope this work demonstrates a productive synthesis of these perspectives, and takes a step forward in the development of more intelligent autonomous systems.
Acknowledgments
This work is supported by ONR MURI #N00014-22-1-2773, ONR #N00014-21-1-2758, and the National Science Foundation under NSF Award 1922658. AGC is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), PGSD3-559278-2021. Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), PGSD3-559278-2021. We are grateful for insightful discussions with Rajesh Ranganath. We thank Doina Precup, Arun Ahuja, Ben Evans, Siddhant Haldar, Nikhil Bhattasali, Ulyana Piterbarg, Dongyan Lin, and Mandana Samiei for their helpful feedback on earlier drafts of this paper. Finally, we thank the anonymous reviewers for making this work stronger in many ways.
@@ B.1 Abstracted MDP
Here we re-iterate our framework and discuss it in greater details. We first define a low level (reward-free) Markov Decision Process (MDP, Puterman (1994)) as the tuple $\langle\mathcal{S},\mathcal{A},P\rangle$, with (low level) state space $\mathcal{S}$, primitive action space $\mathcal{A}$, and transition probability function $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$. Suppose the agent interacts with the environments at discrete time-steps $u=0,1,2,...$, then $P(s_{u+1}|s_{u},a_{u})$ describes the probability of transitioning to state $s_{u+1}\in\mathcal{S}$ after one time-step when choosing action $a_{u}\in\mathcal{A}$ in the current low level state $s_{u}$.
@@
B.1.1 Interpretation via Options Framework
It is natural to interpret our hierarchical framework within the Options framework (Sutton et al., 1999). An option consists of a policy $\pi:\Omega\times\mathcal{A}\rightarrow[0,1]$, a termination function $\beta:\Omega\rightarrow[0,1]$, and an initiation set of states $\mathcal{I}\subseteq\mathcal{S}$. $\Omega$ is used to denote the set of all possible historical states.
@@ B.1.2 A note on semi-MDP
The decision process that selects amongst options is a semi-MDP (Theorem 1, (Sutton et al., 1999)). The general way to model semi-MDP dynamic requires considering the amount of low level timesteps that passes between each abstract level decision as a random variable (Sutton et al., 1999; Barto & Mahadevan, 2003). Instead, we chose to simply treat the abstracted level as having a fixed time interval in between steps. This is chosen for simplicity as our focus is with learning efficiency on the abstract level, and this formulation allow us to directly apply efficient algorithms developed to solve MDPs. We also see empirically that given our $\mathcal{M}$ and $\mathcal{B}$ this choice does not lead to a less optimal policy compared to algorithms that directly solve the low level MDP.
B.1.3 Relationship to other MDP formalism
Our Ab-MDP is related to Factored MDPs (Boutilier et al., 1995; 1999; Degris & Sigaud, 2013). Factored MDPs model states as decomposable into a set of factors $X_{t}=\{x^{(1)},x^{(2)},...\}$, with each factor having domain $\text{Dom}(x^{(i)})$. This allows the transition probability function to be decomposed into Dynamic Bayesian Networks (DBNs) between the factors at two successive time-steps $t-1$ and $t$, with the factors as the nodes of the DBN. Concretely, this decomposes the transition probability function for each action into a product of conditional probabilities,
$P_{a}(X_{t}|X_{t-1})=\prod_{i}P_{a}(x^{(i)}_{t}|\text{Parent}(x^{(i)}_{t}))\,,$ (6)
where $\text{Parent}(x^{(i)}))$ denotes the factors in $X_{t-1}$ which are the parent of $x^{(i)}_{t}$. In the most generic form, each action $a\in\mathcal{A}$ induces a different DBN. If the DBN is known or well-approximated, the Factored MDP can be efficiently solved (Guestrin et al., 2003).
Similarly, the Ab-MDP considers a decomposition of state into items (analogous to factors), and our method MEAD shares the philosophy of efficient utilization of a factorized state space. While one could interpret Ab-MDP as a kind of Factored MDP with discrete domains, the two differs in their emphasis on transition vs. actions. Factored MDPs defines arbitrary actions, each inducing a potentially different conditional independence structure in the transition dynamics. The Ab-MDP instead defines a set of next state changes as the actions, without explicitly specifying any independence structures in its transitions dynamics. Our method MEAD then makes use of this action representation to directly model the probability of item-attribute changes, while implicitly learning the transition dynamics in the weights of the neural network.
Our Ab-MDP is also related to the Object-oriented MDP (OO-MDP) of Diuk et al. (2008), and shares a similar structure of describing a set of “objects” and their attributes separately. In OO-MDP, an environment consists of a set of objects $O=\{o_{1},...,o_{o}\}$, each belonging to one of $c$ classes $\{C_{1},...,C_{c}\}$, and each having a set of (class) attributes $\text{Att}(C)=\{C.{a_{1}},...,C.{a_{a}}\}$. An OO-MDP “object $o.\text{state}$—is a value assignment to all of its attributes ($\{a_{1},...,a_{a}\}$), and the OO-MDP “state” is the union of all of its object states: $s=\cup_{i=1}^{o}o_{i}.\text{state}$.
Our Ab-MDP similarly contains a set of items, $X=\{x^{(1)},...,x^{(N)}\}$, with each item consisting of an item identity and an attribute, $x^{(i)}=(\alpha^{(i)},\xi^{(i)})$. We can view our Ab-MDP as a soft generalization of OO-MDP by modelling:
  • The OO-MDP object (and its associated class) as an Ab-MDP item identity: $\alpha^{(i)}=o_{i}$,
  • The OO-MDP attribute as the Ab-MDP item attribute: $\xi^{(i)}=\{C.{a_{1}},...,C.{a_{a}}\}$, $o_{i}\in C$.
To model state transitions, OO-MDP makes additional assumptions that objects interact through their class identities. Specifically, a set of relations determine if two objects interact, with resulting effects that is used to update the object state. This requires the designer to additionally provide a set of accurate relations functions and effects, with the benefit of facilitating within-class generalization. In contrast, Ab-MDP makes no additional assumption about object class, relations, or interactions between objects. We instead directly models (probabilistically) whether an effect is possible. This is done in a fully data-driven way through a neural function $f:\mathcal{X}\times\mathcal{B}\rightarrow[0,1]$ and empirically is shown to generalize well across item-attribute variations. Item interactions are implicitly represented by the attention mechanism over the set of input items.
The neural constraint satisfaction (NCS) approach of Chang et al. (2023) introduces an entity-set state representation, where each entity $h^{(i)}$ is decomposable into type $z^{(i)}$ and entity state $s^{(i)}$. Our Ab-MDP can naturally be mapped onto their entity-state abstraction by setting an Ab-MDP item identity as the NCS type ($\alpha^{(i)}=z^{(i)}$) and the Ab-MDP item attribute as the NCS entity state ($\xi^{(i)}=s^{(i)}$). Note the Ab-MDP “behaviour” (i.e. abstract action) is equivalent to the NCS “action”.
NCS makes additional assumptions that attribute change is dependent only the behaviour: $\xi^{(i)}_{t+1}=f(\xi^{(i)}_{t},b_{t})$. On the other hand, Ab-MDP models more general relationships $\xi^{(i)}_{t+1}=f(\alpha^{(1)}_{t},\xi^{(1)}_{t},...,\alpha^{(N)}_{t},\xi^{(N)}_{t},b_{t})$ (e.g. Equation 1). The NCS assumption is restrictive in that it would fail to capture cases where an item’s attribute change depends on what the item’s identity is, and/or the attributes of other items.
The Deep Latent Particles (DLP) approach used in Haramati et al. (2024) learns a latent space containing a set of $K$ particles: $z=\{(z_{p},z_{s},z_{d},z_{t},z_{f})_{i}\}_{i=1}^{K-1}$. $z_{p}\in\mathbb{R}^{2}$ describes the particle’s (x,y) position on an image, $z_{s}\in\mathbb{R}^{2}$ describes its (x,y) scale, $z_{d}\in\mathbb{R}$ describe its “depth”, $z_{t}\in\mathbb{R}$ its transparency, and $z_{f}\in\mathbb{R}^{l}$ the visual features around the particle. We can cast this into the Ab-MDP by having $K$ items whose item identities is the particle’s index, $\alpha^{(i)}=i$, and item attribute is a $6+l$-dimensional vector describing the particle features; $\xi^{(i)}=(z_{p},z_{s},z_{d},z_{t},z_{f})_{i}$. DLP can be one way of getting abstraction from pixels, albeit it is restricted to settings where the state is image-based.
B.1.4 Example applications
We provide further examples of possible Ab-MDP constructions for a few common settings. Note we provide one example abstraction per domain, though multiple types of Ab-MDP abstractions can in theory be constructed depending on the choice of the item set and attribute set.
Environment Description Example abstract states How to get abstract behaviours
Kitchen (Gupta et al., 2019) A robot arm is in a MuJoCo kitchen with multiple things to do such as turning on lights, turning on kettle, opening oven / microwave / cabinet, etc. Item set: {light, oven, microwave, cabinet}. Attribute set: {open, closed} Imitating expert play data in Gupta et al. (2019) which already produce trajectories that lead to abstract state changes
Cube pushing (Haramati et al., 2024) A robot arm pushes cubes on a table to their desired positions Item set: {red cube, blue cube, green cube, …}. Attribute set: {at goal position, not at goal position} RL to push a single block to goal at a time
Real world robot (Lee et al., 2024) A Stretch robot in a kitchen, moving objects in and out of containers Item set: {drawer, bag, oven, box, bread, can}. Attribute set: {open, closed, on table, in bag, in drawer, in oven} Human imitation learning data containing trajectories between abstract state changes
Table 1: Example Ab-MDP’s applied to various domains
B.1.5 Multi-Item Behaviours
@@ B.2.1 The generality of Ab-MDP
We make the observation that any sparse reward RL task can be described by a Ab-MDP with a single item,
@@ C.1 Model Fitting
Appendix C.2 contains architectural details for $f_{\theta}$. All in all, we found that our sampling method and loss function lead to a stable, simple-to-optimize objective. Moreover, it is possible to directly use $\rho$ to do planning, as is done in Zhang et al. (2018). However, we show in Section F.1.3 that it is more efficient to use our parametric modelling approach as it generalizes better to new objects.
@@ C.2 Model Architecture
We use a small key-query attention architecture that takes in sets of vectors to produce a binary prediction (Figure 11). Each item $x^{(i)}=(\alpha^{(i)},\xi^{(i)})$ is represented by its item identity and item attribute embeddings. An Ab-MDP state containing a set of $m$ items is represented as a $m\times(d_{\text{iden}}+d_{\text{attr}})$ matrix. The behaviour $b=(\alpha^{(i)},\xi^{\prime})$ describes an item identity and the new attribute to try to change it to, and is a $d_{\text{iden}}+d_{\text{attr}}$ dimensional vector. The set of items and the behaviour is passed element-wise through a small MLP (specifically a 2 layer ReLU network with 128 units in its hidden layer) to produce a set of $m\times d_{\text{emb}}$ item embeddings, and a $d_{\text{emb}}$-dimensional behaviour embedding ($d_{\text{emb}}=128$). The item embeddings are linearly projected into key and value matrices $K$ and $V$ both with shape $m\times d_{k}$, and the behaviour embedding is linearly projected into a $1\times d_{k}$ query vector $Q$ ($d_{k}=256$). We compute the scaled dot product attention (Vaswani, 2017):
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}V\right)\,.$ (11)
The output of the attention operation is an $1\times d_{k}$ vector. Finally, it is linearly projected to a scalar, representing the un-normalized logit for binary prediction.
@@ C.3 Exploration
Intuitively, a state with many visits (large $\text{N}(X,b)$) will have lower reward, while $T$ prevents the intrinsic reward from shrinking to zero (so the agent explores indefinitely). Indeed, in the bandits setting, this reward function is maximized when all states are visited uniformly (Zhang et al., 2018). We emphasize that all of our model training is done with this intrinsic reward, with no task-specific information injected at training time.
@@ Appendix D Environments
Figure 12: Example episode in an example 2D Crafting environment, figure taken from Chen et al. (2020). To solve this task, the agent needs to pickup the pickaxe, mine the cobblestone stash to get cobblestone, then use the acquired cobblestone to craft cobblestone stairs. The items must be acquired in that order as each step requires having the item from the previous step (with an additional obstacle of being able to cross the door only if it has a key).
We adopt three crafting environments proposed in Chen et al. (2020), which contain Minecraft-like crafting logic. The agent needs to traverse a grid world, collect resources, and build increasingly more complex items by collecting and crafting ingredients (e.g. getting an axe, chopping tree to get wood, before being able to make wooden planks and furniture that depend on planks). Primitive actions consist of movement in the four cardinal directions (N,E,S,W), toggle (of a switch to open doors), grab (e.g. pickaxe), mine (e.g. ores), craft (at crafting bench of specific items). The low level observation space consist of a dictionary containing: (i) an (5, 5, 300) matrix describing a $5\times 5$ world with each state being described by an 300-dim word embedding, (ii) an 300-dim inventory embedding describing things in the inventory, and (iii) a 300-dim goal embedding describing in words the task to complete. We describe each environment’s crafting sequences required to finish each task in Table 2.
@@ D.2 MiniHack
We also use the MiniHack environment (Samvelyan et al., 2021), which is built on the Nethack Learning Environment (Küttler et al., 2020). We use five hard-exploration games from Henaff et al. (2022) requiring interaction with different types of objects. The base game has a multi-modal observation space involving image, text, inventory, etc. Solving the game from the base observation is extremely difficult since each environment has multiple variants of an object (e.g. Freeze-Wand has 27 different types of wands), and randomly generated object locations. Indeed, Henaff et al. (2022) found that performant RL methods such as IMPALA (Espeholt et al., 2018; Küttler et al., 2019) and all tested global exploration methods (ICM (Pathak et al., 2017) and RND (Burda et al., 2019)) fail to solve these environments.
@@
We then embed the object identities and attributes before providing them to the agents. For object identities, we take each object’s glyph number, and map it to the NetHack text description for that glyph (which we can extract from the screen_descriptions modality in the low level observation). For each glyph’s text description (which is a short sentence), we pre-process the sentence to remove stop words, then take each word’s GloVe (Pennington et al., 2014) embedding and averaging over the sentence. We use the 300-dimensional Wikipedia 2014 + Gigaword 5 GloVe word vector embeddings from https://nlp.stanford.edu/projects/glove/ (6B tokens, 400k vocab, uncased). The 300-dim vectors were chosen as we found them to have naturally good separation between object types in preliminary analysis.
@@ D.2.1 Sandbox Environment
(a) Levitate-Potion
(b) Freeze-Random
(c) Sandbox
Figure 14: Example Minihack environments. Only a cropped version of the screen is rendered for presentation clarity. The full observation space contains a larger screen, agent’s inventory, in-game stats, and in-game messages (not shown here).
We design a sandbox environment where the agent can observe and interact with all items it may encounter in one of the five evaluation environments. In the sandbox, each of these objects has an 0.5 chance (independently) of spawning at a random location: levitate boots, levitate rings, levitate potions, freeze horn, freeze wands. Further there is a 0.5 change that a second of the same object type will spawn. The agent can interact with any of the objects for up to 250 low level frames, or the episode terminates once the agent sets any of the object’s attribute to “used”. Figure 14(c) shows an instantiation of an episode.
@@ D.2.2 Low Level Policy Errors
    To use an item in the inventory, the policy needs to select which item to use (e.g. to drink a potion the policy needs to first do primitive action ”quaff”, then primitive action that selects the inventory object to drink, etc.). When there are multiple objects in the inventory it just picks a random object, meaning it might not pick the object specified by the behaviour
@@
    The “go stand on” policy is coded to move in a straight line towards an object. Since we measure a policy ”success” whenever the abstract state changes, if the agent walks over another object en route, the Ab-MDP will register an abstract state change and deem the transition unsuccessful
@@
    To successfully cast a freeze spell, the NetHack game needs to return a message saying the spell has bounced off a wall. To our knowledge, this is stochastic—we got as far as figuring out that the spell needs to travel sufficient distance to bounce, but even so we were unable to produce a low level policy that results in spell bounce 100% of the time. This succeeds $\sim$ 80% of the time